OpenStudy (meilendurcer):

–4 1/2• –2 1/3= A. 10 1/2 B. -10 1/2 C. -9 D. 9

jhonyy9 (jhonyy9):

so little help ,if 2 1/3 = (3*2 +1)/3 =(6+1)/3 = 7/3 so can you calcule hence your exercise too ?

OpenStudy (meilendurcer):

i can try

jhonyy9 (jhonyy9):

ok the first i will calcule it so -4 1/2 = -(4*2 +1)/2 = -(8+1)/2 = -9/2

jhonyy9 (jhonyy9):

can you calcule the same -2 1/3 =

OpenStudy (meilendurcer):

3?

jhonyy9 (jhonyy9):

no the same like how i have calculet it -4 1/2 ok ?

OpenStudy (meilendurcer):

ok how do i do that?

jhonyy9 (jhonyy9):

step by step how i have doit -4 1/2

jhonyy9 (jhonyy9):

so there are - 2 1/3 = - (2*3 +1)/3 =

jhonyy9 (jhonyy9):

courage

jhonyy9 (jhonyy9):

2*3 =

OpenStudy (meilendurcer):

6

jhonyy9 (jhonyy9):

6+1=

OpenStudy (meilendurcer):

7

jhonyy9 (jhonyy9):

-(2*3 +1) =

jhonyy9 (jhonyy9):

-(6+1) =

jhonyy9 (jhonyy9):

-7

jhonyy9 (jhonyy9):

so there will be -7/3

jhonyy9 (jhonyy9):

than your exercise will be after these calcules how ?

OpenStudy (meilendurcer):

ok

jhonyy9 (jhonyy9):

can you rewrite it here now please ?

OpenStudy (meilendurcer):

yes

jhonyy9 (jhonyy9):

can i look it now here ?

OpenStudy (meilendurcer):

\[-\frac{ 7 }{ 3 } \div = 2.3 ?\]

jhonyy9 (jhonyy9):

how you have got -7/3 ?

OpenStudy (meilendurcer):

oh ok i understand 1 sec

jhonyy9 (jhonyy9):

-9/2 * -7/3 =

jhonyy9 (jhonyy9):

so you need multiplie these fractions ?

OpenStudy (meilendurcer):

okay

jhonyy9 (jhonyy9):

so than what will be the right answer ?

OpenStudy (meilendurcer):

10.32?

jhonyy9 (jhonyy9):

do you see this there ?

OpenStudy (meilendurcer):

no

jhonyy9 (jhonyy9):

so than how do you think ,is right ?

OpenStudy (meilendurcer):

i don't i just don't know how to do this type of math..

jhonyy9 (jhonyy9):

how multiplie two fractions ?

jhonyy9 (jhonyy9):

a/b * c/d =

OpenStudy (meilendurcer):

yea

jhonyy9 (jhonyy9):

so a c --- * --- = b d

jhonyy9 (jhonyy9):

so this is equal (a*c)/(b*d) ok ?

jhonyy9 (jhonyy9):

can you calcule your exercise the same ?

OpenStudy (meilendurcer):

i got it. my answer is d

jhonyy9 (jhonyy9):

so how you have got it ? not is right can you calcule it here ?

jhonyy9 (jhonyy9):

-9/2 * -7/3 = (-9 * -7)/(2*3) = ok ? so than this will be equal ?

jhonyy9 (jhonyy9):

minus time minus equal

OpenStudy (meilendurcer):

ok

jhonyy9 (jhonyy9):

courage please

jhonyy9 (jhonyy9):

so than how many will be ?

OpenStudy (meilendurcer):

378?

jhonyy9 (jhonyy9):

-9 -7 --- * ----- = ------ 2 3

OpenStudy (meilendurcer):

-27 and -14?

jhonyy9 (jhonyy9):

why ? -9 *-7 ------ = 2*3

OpenStudy (meilendurcer):

63 and 6?

jhonyy9 (jhonyy9):

yes 63/6 =

OpenStudy (meilendurcer):

10.5

jhonyy9 (jhonyy9):

can you prove me here that a. is the right answer

jhonyy9 (jhonyy9):

???

OpenStudy (meilendurcer):

isn't 10 .5 mean 10 and 1/2?

jhonyy9 (jhonyy9):

so there is 63/6 yes ?

OpenStudy (meilendurcer):

yes

jhonyy9 (jhonyy9):

63 divided by 6 = 10 and remainder 3/6 but you know that 3/6 equal 1/2 yes ? so than this result 10 1/2 ok ?

OpenStudy (meilendurcer):

ok thank you

jhonyy9 (jhonyy9):

was easy ??? your welcom good luck bye